OpenStudy (brynndelyn):

A toy company is considering a cube or a sphere shaped container for packaging a new product. The height of the cube would equal the diameter of the sphere. Compare the volume-to-surface area ratios of the containers. Which packaging will be more efficient? For a sphere, SA=4(pi)r^2.

OpenStudy (brynndelyn):

@dan815 @rishavraj

rishavraj (rishavraj):

for a cube length = width = height\ let the height be = x so volume of cube = x^3 and surface area = 6 x^2

rishavraj (rishavraj):

height of cube = diameter of sphere so diameter = x radius of sphere = x/2

rishavraj (rishavraj):

@brynndelyn done???????